OpenStudy (anonymous):

Find f '(a). f(t) = 6t + 38/t + 7 Please show step by step..thanks

OpenStudy (anonymous):

Okay :) Well I would first rewrite this into a form that seems familiar and easy to work with. Rewrite this as f(t)=6t+38(t^(-1))+7. From here, apply the power rule to find the derivative. f'(t)=6+38(-1)(t^(-2)). f'(t)=6-38/t^2. From here you can just plug in your a value for t. Rendering: f'(a)=6-38/a^2. Two points of note if you are confused. The constant(7) goes to zero when you differentiate it. The second is that the power rule says: \[d/dx(x^n)=(n)x^{n-1}\]

myininaya (myininaya):

(6t)'=6 (38/t)'=(38t^{-1})=-1*38t^{-1-1}=-38t^{-2}=-38/t^2 (7)'=0

myininaya (myininaya):

f'(t)=6-38/t^2 f'(a)=6-36/a^2